How would I solve this question?
A high speed train travels from Chicago to Denver at an average speed of 90 miles per
hour, but can make the return trip at an average speed of 110 miles per hour. What is the
average speed for the round trip?

Answer by Alan3354(69443)

A high speed train travels from Chicago to Denver at an average speed of 90 miles per
hour, but can make the return trip at an average speed of 110 miles per hour. What is the
average speed for the round trip?
------------
d = distance one way
t = RT time = d/90 + d/110
RT distance = 2d
----
RT avg speed = 2d/(d/90 + d/110)
= 2/(1/90 + 1/110) = 2/((90 + 110)/(90*110))
= 2*90*110/(90 + 110) = 99 mi/hr
----------------------
Avg speed for a round trip = 2*r1*r2/(r1 + r2)
Distance is not a factor.